Solve the initial value problem (IVP) Ut + 3ux + 3u 0, u(x,0) = x2, (x, t) ER [0, +00).
Answer #1

now et tsüxt su = with U (20) = x² (x, t) t RX [0, A) Lagrange Sub ecuation is da du Subsidiary dt در -3ce 3 ratio now of coyCY occu) Ø (x-3t) uex := U (^,t) - e-X (9-36) srence U (2,0) = forum (a) u Carol e-2 Ø ca] x2 as u 12,0) = x² = n² = era e-2